A line passes through the point (-7,4) and has a slope of -2 . Write an equation for this line.

Answer by Cromlix(4381)

Using equation:
y - b = m(x - a)
a = -7, b = 4 and slope (m) = -2
y - 4 = -2(x - (-7))
y - 4 = -2(x + 7)
y - 4 = -2x - 14
y = -2x - 14 + 4
y = -2x - 10
Hope this helps.
:-)